# MedGellan: LLM-Generated Medical Guidance to Support Physicians in Diagnosis

### v2

@@ 1 Introduction
Based on these insights, we propose MedGellan, a novel pipeline where an assistant LLM is employed to analyze and provide guidance for diagnosing a patient’s health condition and in the later stage, a doctor, instead of studying the raw data or the electronic health record (EHR), can take advantage of the guidance in order to make the final diagnosis (see Figure 1-left).
@@ 3 MedGellan
MedGellan is designed to assist physicians in predicting accurate discharge diagnoses. An overview of the proposed framework is shown in Fig. 1-right. In brief, MedGellan consists of two modules. The first module is powered by the assistant LLM —a state-of-the-art LLM—which takes the triage note and radiology report as input to generate clinical guidance. We ask the assistant LLM to observe the triage information at first. Next, the assistant LLM is asked to look at the radiological report and update its assessment. Finally, based on its updated information, it generates a comprehensive guidance on the patient’s health condition that would help the physician to determine the ideal diagnosis. The second module is dedicated to the physician. The physician uses the guidance obtained from the first module and makes diagnosis for the patient. During this phase, the physician is allowed to examine only the guidance, unlike the first module where the assistant LLM gets access to the raw information, i.e., the triage and the radiology report.
@@ 4 Experimental Work
We combine MIMIC-CXR (Johnson et al., 2019), MIMIC-IV-ED (Johnson et al., 2020) and MIMIC-IV (Johnson et al., 2024) datasets for the chest radiographs, triage information, and diagnosis, respectively. Table 1 summarizes the key components of each dataset. Mimic-IV is a comprehensive dataset containing health records of all patients who have visited the hospital. Mimic-IV-ED is a subset focused specifically on emergency department data, while MIMIC-CXR provides an extensive collection of chest radiograph images.
@@
As outlined in Section 3, the first task involves generating clinical guidance using an LLM, which is then reviewed by a physician. While the framework is designed to include a human physician, for preliminary experiments aimed at assessing its potential, we simulate the physician’s role with an LLM. We use Llama 3–70B (Grattafiori et al., 2024) as our assistant LLM model. For the physician LLM model, we experiment with Llama 3 (8B and 70B) (Grattafiori et al., 2024), Gemma 2–27B (Team et al., 2024), and Qwen2–72B (Team, 2024). All experiments are conducted using the Ollama framework. [^2] https://github.com/ollama.
@@
Prompt for generating Guidance(Bayes strategy) You are an expert physician assistant trained to analyze patient records and generate a structured, evidence-weighted summary to aid in diagnosis. Your role is to synthesize information probabilistically, emphasizing prior observations (triage data) and new evidence (radiology findings) to refine the clinical understanding of the case.
### Input Data:
@@ 4.1 Results
In Table 2, we compare the performance of the physician LLM across three input settings: (i) using only the triage note, (ii) using both the triage note and the radiology report, and (iii) using the full MedGellan framework, which includes intermediate guidance. Predicting multiple ICD-10 codes for a single patient constitutes a challenging multi-label classification task. We evaluate performance using precision, recall, and $F_{1}$ score. Across all physician LLM variants used to simulate a real physician, the guidance provided by the MedGellan framework consistently outperforms the other two baselines in terms of recall and $F_{1}$. While we observe a slight drop in precision when guidance is included, the gains in recall and $F_{1}$—metrics more critical in high-stakes medical decision-making—indicate a favorable trade-off. This suggests that incorporating proper guidance enables the model to be more comprehensive in its predictions, thereby reducing the risk of false negatives.
[DB: added non bayes]

### v3

@@ 1 Introduction
Based on these insights, we propose MedGellan, a novel pipeline where an assistant LLM is employed to analyze and provide guidance for diagnosing a patient’s health condition and in the later stage, a doctor, instead of studying the raw data or the electronic health record (EHR), can take advantage of the guidance in order to make the final diagnosis (see Figure˜1-left).
@@ 3 MedGellan
MedGellan is designed to assist physicians in predicting accurate discharge diagnoses. An overview of the proposed framework is shown in Fig.˜1-right. In brief, MedGellan consists of two modules. The first module is powered by the assistant LLM —a state-of-the-art LLM—which takes the triage note and radiology report as input to generate clinical guidance. We ask the assistant LLM to observe the triage information at first. Next, the assistant LLM is asked to look at the radiological report and update its assessment. Finally, based on its updated information, it generates a comprehensive guidance on the patient’s health condition that would help the physician to determine the ideal diagnosis. The second module is dedicated to the physician. The physician uses the guidance obtained from the first module and makes diagnosis for the patient. During this phase, the physician is allowed to examine only the guidance, unlike the first module where the assistant LLM gets access to the raw information, i.e., the triage and the radiology report.
@@ 4 Experimental Work
We combine MIMIC-CXR (Johnson et al., 2019), MIMIC-IV-ED (Johnson et al., 2020) and MIMIC-IV (Johnson et al., 2024) datasets for the chest radiographs, triage information, and diagnosis, respectively. Table˜1 summarizes the key components of each dataset. Mimic-IV is a comprehensive dataset containing health records of all patients who have visited the hospital. Mimic-IV-ED is a subset focused specifically on emergency department data, while MIMIC-CXR provides an extensive collection of chest radiograph images.
@@
As outlined in Section˜3, the first task involves generating clinical guidance using an LLM, which is then reviewed by a physician. While the framework is designed to include a human physician, for preliminary experiments aimed at assessing its potential, we simulate the physician’s role with an LLM. We use Llama 3–70B (Grattafiori et al., 2024) as our assistant LLM model. For the physician LLM model, we experiment with Llama 3 (8B and 70B) (Grattafiori et al., 2024), Gemma 2–27B (Team et al., 2024), and Qwen2–72B (Team, 2024). All experiments are conducted using the Ollama framework. [^2] https://github.com/ollama.
@@
Prompt for generating Guidance You are an expert physician assistant trained to analyze patient records and generate a structured, evidence-weighted summary to aid in diagnosis. Your role is to synthesize information probabilistically, emphasizing prior observations (triage data) and new evidence (radiology findings) to refine the clinical understanding of the case.
### In
put Data:
@@ 4.1 Results
In Table˜2, we compare the performance of the physician LLM across three input settings: (i) using only the triage note, (ii) using both the triage note and the radiology report, and (iii) using the full MedGellan framework, which includes intermediate guidance. Predicting multiple ICD-10 codes for a single patient constitutes a challenging multi-label classification task. We evaluate performance using precision, recall, and $F_{1}$ score. Across all physician LLM variants used to simulate a real physician, the guidance provided by the MedGellan framework consistently outperforms the other two baselines in terms of recall and $F_{1}$. While we observe a slight drop in precision when guidance is included, the gains in recall and $F_{1}$—metrics more critical in high-stakes medical decision-making—indicate a favorable trade-off. This suggests that incorporating proper guidance enables the model to be more comprehensive in its predictions, thereby reducing the risk of false negatives.